# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v2

@@ 3.4 Post-processing
Once the series of output probabilities $P=(p_{i})_{i=1}^{N}$ for the entire video $V$ are obtained, we perform MaE spotting and ME spotting independently, following the methodology outlined in [leng2022abpn]. We will elaborate on the ME spotting process, with the MaE spotting process following a similar method. Initially, we identify all potential ME apex frames $U^{\mathrm{me},\mathrm{apex}}$ based on the criterion $p_{l}^{\mathrm{me},\mathrm{apex}}>\theta_{\mathrm{apex}}$, where $\theta_{\mathrm{apex}}$ is a threshold. For each spotted apex frame $u^{\mathrm{me},\mathrm{apex}}_{l}\in U^{\mathrm{me},\mathrm{apex}}$ with the frame index $l$, we select the onset frame with the highest onset probability from the left side of the apex frame within the range of $l-\frac{k^{\mathrm{me}}}{2},l-\frac{j^{\mathrm{me}}}{2}]$. Similarly, we select the offset frame with the highest offset probability from the right side of the apex frame within the range of $l+\frac{j^{\mathrm{me}}}{2},l+\frac{k^{\mathrm{me}}}{2}]$, where $k^{\mathrm{me}}$ and $j^{\mathrm{me}}$ represent the average duration and minimum duration of a ME, respectively. As a result, a ME proposal $\phi_{l}$ is obtained, including the onset frame, offset frame, and expression type. Subsequently, we assign a score $s_{l}=p_{b}^{\mathrm{me},\mathrm{onset}}\times p_{l}^{\mathrm{me},\mathrm{apex}}\times p_{d}^{\mathrm{me},\mathrm{offset}}$ to $\phi_{l}$, where $b$ and $d$ denote the frame indices of the onset frame and offset frame selected by the aforementioned rule, respectively.
@@ 4.1 Experimental setup
Datasets. We conduct experiments on two datasets: SAMM-LV [yap2020samm], CAS(ME)2 [qu2017cas], and CAS(ME)3 [9774929]. The SAMM-LV dataset includes 147 raw long videos with 343 MaE clips and 159 ME clips. The CAS(ME)2 dataset includes 87 raw long videos with 300 MaE clips and 57 ME clips. The large-scale dataset CAS(ME)3 includes 956 raw long videos with 3490 MaE clips and 1109 ME clips. Since the frame rate of the SAMM-LV dataset is 200fps while the frame rate of the CAS(ME)2 and CAS(ME)3 datasets is 30fps, we subsample every 7th frame from the SAMM dataset to align the frame rates of all three datasets.
Evaluation metric. Following the protocol of MEGC2021 [li2021fme], we evaluate our model on each dataset separately and employ a leave-one-subject-out cross-validation strategy in our experiments. The true positive (TP) is defined based on the Intersection over Union (IoU) between the proposal and the ground-truth expression clip. Specifically, given a ground-truth expression clip and its expression type, we compare it with all expression proposals that have the same estimated expression type. An expression proposal $W_{\mathrm{Proposal}}$, is considered a TP when it satisfies the following condition:
@@
Training details. We use the AdamW optimizer [loshchilov2017decoupled] to optimize our model, setting the learning rate to 0.0002, $\beta_{1}$ to 0.7, and $\beta_{2}$ to 0.9. The temperature parameter $\tau$ in Eq. (4) is set to 0.5. We train our model for 100 epochs with a batch size of 512. The facial images input for the SW-MRO feature extraction are resized to $(320,320)$. The general expression durations, $k^{\mathrm{mae}}$ and $k^{\mathrm{me}}$, are set to 60 and 16, respectively, while the minimum expression durations, $j^{\mathrm{mae}}$ and $j^{\mathrm{me}}$, are set to 16 and 4. The hidden dimension and the number of heads in the MHSA mechanism is set to 256 and 4, respectively.
@@ 4.2 Comparison with state-of-the-art methods
The remarkable improvements in ME spotting are mainly attributed to our proposed SW-MRO feature. Previous methods generally extract optical flow between adjacent frames and employ either a pretrained I3D model or simple neural networks to model these coarse motion patterns. In contrast, our proposed SW-MRO feature effectively extracts subtle motions while avoiding the optical flow being dominated by head movements. Moreover, the carefully designed SpotFormer can comprehensively analyze the latent motion patterns within the SW-MRO feature. In addition, the proposed supervised contrastive learning further enhances discriminative feature learning between different expression types, leading to improved model performance.
We further validate our method on the CAS(ME)3 dataset [9774929] and the results are shown in Table II. Following LGSNet [yu2023lgsnet], we pre-process the dataset by considering only ground-truth MEs with a duration of less than 15 frames and ground-truth MaEs with a duration between 16 and 120 frames for training and validation, ensuring a fair comparison. This data pre-processing method results in 2231 ground-truth MaEs and 285 ground-truth MEs. The results demonstrate that our method outperforms LGSNet [yu2023lgsnet], especially in ME spotting, highlighting the strong generalization ability of our method.
@@ 4.3 Ablation studies
We conduct ablation studies to evaluate our proposed modules and Table III shows the experimental results. The acronyms SW-MRO and FLGP denote specific modules in our model. Specifically, SW-MRO involves calculating sliding window-based multi-temporal-resolutionoptical flow features instead of optical flow between adjacent frames, while FLGP integrates facial local graph pooling operation into SpotFormer for multi-scale spatial feature learning instead of global average pooling over facial graphs. The results show the effectiveness of each proposed module. Notably, the introduction of the SW-MRO enhances the extracted motion features, particularly in revealing subtle motions that exist in MEs. This enhancement leads to a significant overall performance improvement of 17.0%/17.7% on the SAMM-LV and CAS(ME)2 datasets. Especially in ME spotting, it has led to an improvement of 69.2%/200.1% on the SAMM-LV and CAS(ME)2 datasets. The introduction of the FLGP operation for multi-scale spatial feature learning results in a further improvement of 3.0%/3.4% on the SAMM-LV and CAS(ME)2 datasets compared to using global average pooling. This demonstrates the superior representational capabilities of our proposed model.
Figure 5: Some visualization optical flow of certain micro-expression frames computed by three strategies. The data comes from the vertical component of optical flow computed at the left mouth corner when subject 11 from the SAMM-LV dataset is performing a micro-expression. (a) optical flow computed between adjacent frames; (b) optical flow computed using the proposed SW-MRO; (c) optical flow computed with a large sliding window strategy.
@@
SW-MRO FLGP MaE ME Overall MaE ME Overall
@@
Fig. 5 shows the qualitative comparison between different optical flow extraction strategies. When computing optical flow between adjacent frames, the motions present in MEs are so subtle that the noise in the optical flow could overshadow these subtle movements, impacting the quality of the data and making it difficult to reveal these subtle expressions. On the other hand, employing a large sliding window strategy for optical flow computation introduces substantial interference from irrelevant motions such as head movements, posing challenges in accurately detecting facial expressions. On the contrary, the utilization of the SW-MRO helps alleviate these problems. It achieves a balance between amplifying subtle facial muscle motions and mitigating the impact of head movements.
@@ 4.4.1 Multiple model architectures
TABLE IV: Comparison of multiple model architectures on SAMM-LV and CAS(ME)2.
@@
We validate and compare our proposed multiple multi-scale spatio-temporal model architectures on the SAMM and CAS(ME)2 datasets, and the experimental results are shown in Table IV. With the exception of Paral-ST, our models use the same number of parameters for a fair comparison.
@@
4.4.2 Temporal window size
We further explore how much temporal information is needed to accurately evaluate one frame. Specifically, we compare different temporal window sizes (i.e., temporal receptive field) for the SW-MRO on the SAMM-LV dataset, and the experimental results are presented in Table V. The results indicate that the optimal performance is obtained when the temporal window size is set to 17, which demonstrates the effectiveness of our core idea: perceiving complete MEs and distinguishing between MaEs and MEs. Specifically, the temporal boundary for distinguishing MaEs and MEs is 0.5 seconds, which corresponds to 15 frames when the frame rate is set to 30 fps. A receptive field of 17 frames proves sufficient to perceive a complete ME and amplify the motion information that exists in MEs. When the model can accurately spot MEs, windows that exhibit greater intensity of motion, or windows that exhibit a continuous expression and are hard to perceive a trend of this expression occurring or ceasing within a short period, are more likely to contain MaEs.
TABLE V: Comparison of different temporal window sizes (receptive fields) on SAMM-LV.
Window Size / MaE ME F1-Score
@@ 4.4.2 Temporal window size
TABLE VI: Results of the ablation study on supervised contrastive learning and the choice of hyperparameter.
@@
Figure 7: Visualization analysis of supervised contrastive learning. (a) and (b) show the PCA distribution of certain macro- and micro-expression frames: (a) without supervised contrastive learning and (b) with supervised contrastive learning. (c) and (d) depict the PCA distribution of certain micro-expression frames and neutral frames: (c) without supervised contrastive learning and (d) with supervised contrastive learning.
@@
TABLE VII: Detailed spotting results of the proposed method on SAMM-LV and CAS(ME)2.
@@
4.4.3 Supervised contrastive loss and hyperparameter $\lambda$
Table VI shows the experimental results of introducing supervised contrastive learning into our model and the impact of the hyperparameter $\lambda$ in Eq. (5), which is set to balance classification and supervised contrastive learning. We observe that the introduction of the supervised contrastive loss enables our model to better recognize the classification boundaries between different types of expressions, resulting in an improvement of 6.0%/16.3% on the SAMM-LV and CAS(ME)2 datasets. The optimal value for $\lambda$ is 0.005, increasing $\lambda$ beyond this value starts to impact standard classification, leading to a decrease in spotting accuracy.
Fig. 7 shows the qualitative comparison between scenarios with and without supervised contrastive learning. In the inference phase, we randomly sampled frames and applied principal component analysis (PCA) [bro2014principal] to analyze the distribution of various expression classes. Subsequently, we marked each frame with its corresponding ground-truth expression label. In Fig. 7 (a) and (b), we examine the PCA distribution of specific MaE frames and ME frames with and without supervised contrastive learning. Moving to Fig. 7 (c) and (d), we compare the PCA distribution of specific neutral frames and ME frames with and without supervised contrastive learning. The results indicate that, in the absence of supervised contrastive learning, our model may face challenges distinguishing certain MaE frames from ME frames due to subtle differences. This leads to mixed distributions and misclassification. Additionally, neutral frames may be mistakenly classified as ME frames, as the noise existing in the optical flow introduces slight fluctuations that the model may treat as ME motion information, resulting in incorrect ME proposal detection. With the introduction of supervised contrastive learning into our model, the domain discrepancy increases, thus significantly improving the accuracy of expression spotting.
@@ 4.5 Detailed results
Table VII shows the detailed results on the SAMM-LV and CAS(ME)2 datasets. Our method achieves an overall precision rate of 0.4341/0.4781 and an overall recall rate of 0.4462/0.4902 on SAMM and CAS(ME)2, respectively. Specifically, our method achieves a recall rate of approximately 0.5 in MaE spotting and 0.4 in ME spotting. The exception is for ME spotting on the CAS(ME)2 dataset. The results indicate that it achieves a high precision rate while a low recall rate, which may be due to several factors. One primary reason is the scarcity and imbalance of ME data in the CAS(ME)2 dataset, which contains only 57 ME clips. Some ME samples in CAS(ME)2 involve only eye blinking—a type of motion generally considered irrelevant for facial expression spotting. Furthermore, owing to our model’s strong discriminative capability and the incorporation of supervised contrastive learning, it carefully evaluates ME frames based on limited data. Consequently, it achieves a high precision rate but low recall. In addition, on both datasets, the precision is lower than the recall in MaE spotting, while the opposite is observed in ME spotting. This is mainly due to data imbalance and the specific characteristics of different expression classes in the FES datasets: MaEs are more frequent and generally exhibit higher intensity than MEs. This leads the model to focus more on MaEs and generate higher output probabilities for MaE spotting. As a result, more MaEs—including both TPs and FPs—are detected. In contrast, the model generates lower probabilities for ME spotting, resulting in fewer expression proposals but also fewer FPs.
@@
In summary, our proposed model can achieve a more advanced and balanced performance on both datasets, demonstrating the great generalization ability of our model. In addition, our method requires the detection of facial landmarks and the calculation of optical flows to form a head pose-irrelevant graph, making it more robust to variations such as random head angles. These merits demonstrate the potential of our method for future practical applications.

### v3

@@ 3.4 Post-processing
Once the series of output probabilities $P=(p_{i})_{i=1}^{N}$ for the entire video $V$ are obtained, we perform MaE spotting and ME spotting independently, following the methodology outlined in [27]. We will elaborate on the ME spotting process, with the MaE spotting process following a similar method. Initially, we identify all potential ME apex frames $U^{\mathrm{me},\mathrm{apex}}$ based on the criterion $p_{l}^{\mathrm{me},\mathrm{apex}}>\theta_{\mathrm{apex}}$, where $\theta_{\mathrm{apex}}$ is a threshold. For each spotted apex frame $u^{\mathrm{me},\mathrm{apex}}_{l}\in U^{\mathrm{me},\mathrm{apex}}$ with the frame index $l$, we select the onset frame with the highest onset probability from the left side of the apex frame within the range of $l-\frac{k^{\mathrm{me}}}{2},l-\frac{j^{\mathrm{me}}}{2}]$. Similarly, we select the offset frame with the highest offset probability from the right side of the apex frame within the range of $l+\frac{j^{\mathrm{me}}}{2},l+\frac{k^{\mathrm{me}}}{2}]$, where $k^{\mathrm{me}}$ and $j^{\mathrm{me}}$ define the maximum and minimum temporal ranges for searching boundaries around the apex frame, respectively. As a result, a ME proposal $\phi_{l}$ is obtained, including the onset frame, offset frame, and expression type. Subsequently, we assign a score $s_{l}=p_{b}^{\mathrm{me},\mathrm{onset}}\times p_{l}^{\mathrm{me},\mathrm{apex}}\times p_{d}^{\mathrm{me},\mathrm{offset}}$ to $\phi_{l}$, where $b$ and $d$ denote the frame indices of the onset frame and offset frame selected by the aforementioned rule, respectively.
@@ 4.1 Experimental setup
Datasets. We conduct experiments on three datasets: SAMM-LV [69], CAS(ME)2 [49], and CAS(ME)3 [28]. The SAMM-LV dataset includes 147 raw long videos with 343 MaE clips and 159 ME clips. The CAS(ME)2 dataset includes 87 raw long videos with 300 MaE clips and 57 ME clips. The large-scale dataset CAS(ME)3 includes 956 raw long videos with 3490 MaE clips and 1109 ME clips. Since the frame rate of the SAMM-LV dataset is 200fps while the frame rate of the CAS(ME)2 and CAS(ME)3 datasets is 30fps, we subsample every 7th frame from the SAMM dataset to align the frame rates of all three datasets.
Evaluation metric. Following the protocol of MEGC2021 [31], we evaluate our model on each dataset separately and employ a leave-one-subject-out cross-validation strategy in our experiments. The true positive (TP) is defined based on the Intersection over Union (IoU) between the proposal and the ground-truth expression clip. Specifically, given a ground-truth expression clip and its expression type, we compare it with all expression proposals that have the same estimated expression type. An expression proposal $W_{\mathrm{Proposal}}$, is considered a TP when it satisfies the following condition:
@@
Training details. We use the AdamW optimizer [42] to optimize our model, setting the learning rate to 0.0002, $\beta_{1}$ to 0.7, and $\beta_{2}$ to 0.9. $\theta_{\mathrm{apex}}$ is set to 0.5. The temperature parameter $\tau$ in Eq. (4) is set to 0.5. We train our model for 100 epochs with a batch size of 512. The facial images input for the CSW-MRO feature extraction are resized to $(320,320)$. The maximum temporal ranges for searching boundaries around the apex frame, $k^{\mathrm{mae}}$ and $k^{\mathrm{me}}$, are set to 60 and 16, respectively, while the minimum temporal ranges, $j^{\mathrm{mae}}$ and $j^{\mathrm{me}}$, are set to 16 and 4. The hidden dimension and the number of heads in the MHSA mechanism is set to 256 and 4, respectively.
Figure 5: Some visualization optical flow of certain micro-expression frames computed by three strategies. The data comes from the vertical component of optical flow computed at the left mouth corner when subject 11 from the SAMM-LV dataset is performing a micro-expression. (a) optical flow computed between adjacent frames; (b) optical flow computed using the proposed CSW-MRO; (c) optical flow computed with a large sliding window strategy.
@@ 4.2 Comparison with state-of-the-art methods
The remarkable improvements in ME spotting are mainly attributed to our proposed CSW-MRO feature. Previous methods generally extract optical flow between adjacent frames and employ either a pretrained I3D model or simple neural networks to model these coarse motion patterns. In contrast, our proposed CSW-MRO feature effectively extracts subtle motions while avoiding the optical flow being dominated by head movements. Moreover, the carefully designed SpotFormer can comprehensively analyze the latent motion patterns within the CSW-MRO feature. In addition, the proposed supervised contrastive learning further enhances discriminative feature learning between different expression types, leading to improved model performance.
We further validate our method on the CAS(ME)3 dataset [28] and the results are shown in Table II. Following LGSNet [76], we pre-process the dataset by considering only ground-truth MEs with a duration of less than 15 frames and ground-truth MaEs with a duration between 16 and 120 frames for training and validation, ensuring a fair comparison. This protocol is consistent with those adopted by the SAMM and CAS(ME)2 datasets and also aligns with the general definition that MEs last less than 0.5 seconds, whereas MaEs typically last from 0.5 to 4.0 seconds. The data pre-processing results in 2231 ground-truth MaEs and 285 ground-truth MEs. The results demonstrate that our method outperforms LGSNet [76], especially in ME spotting, highlighting the strong generalization ability of our method.
@@ 4.3 Ablation studies
Proposed CSW-MRO and FLGP. We conduct ablation studies to evaluate our proposed modules and Table III shows the experimental results. The acronyms CSW-MRO and FLGP denote specific modules in our model. Specifically, CSW-MRO involves calculating compact sliding-window-based multi-temporal-resolution optical flow features instead of optical flow between adjacent frames, while FLGP integrates facial local graph pooling operation into SpotFormer for multi-scale spatial feature learning instead of global average pooling over facial graphs. The results show the effectiveness of each proposed module. Notably, the introduction of the CSW-MRO enhances the extracted motion features, particularly in revealing subtle motions that exist in MEs. This enhancement leads to a significant overall performance improvement of 17.0%/17.7% on the SAMM-LV and CAS(ME)2 datasets. Especially in ME spotting, it has led to an improvement of 69.2%/200.1% on the SAMM-LV and CAS(ME)2 datasets. The introduction of the FLGP operation for multi-scale spatial feature learning results in a further improvement of 3.0%/3.4% on the SAMM-LV and CAS(ME)2 datasets compared to using global average pooling. This demonstrates the superior representational capabilities of our proposed model.
@@
CSW-MRO FLGP MaE ME Overall MaE ME Overall
@@
Fig. 5 shows the qualitative comparison between different optical flow extraction strategies. When computing optical flow between adjacent frames, the motions present in MEs are so subtle that the noise in the optical flow could overshadow these subtle movements, impacting the quality of the data and making it difficult to reveal these subtle expressions. On the other hand, employing a large sliding window strategy for optical flow computation introduces substantial interference from irrelevant motions such as head movements, posing challenges in accurately detecting facial expressions. On the contrary, the utilization of the CSW-MRO helps alleviate these problems. These results further confirm that the proposed CSW-MRO feature is not a simple adoption of existing optical flow extraction strategies, but a task-specific design for FES. It achieves a balance between amplifying subtle facial muscle motions and mitigating the impact of head movements. This design is particularly suited for frame-level probability-estimation-based deep learning frameworks, as such models do not require large temporal receptive fields to perceive complete MaEs. Instead, compact windows are sufficient to capture complete MEs while distinguishing between MEs and MaEs. Further discussion is provided in Section 4.4.2.
Figure 6: Comparison of temporal attention maps with and without multi-scale temporal modeling. The predicted frame corresponds to a ground-truth ME apex frame.
Multi-scale temporal modeling. Although 1D CNN-based temporal downsampling has been widely adopted for multi-scale temporal feature extraction [78], we explicitly evaluate its necessity in our framework. Specifically, we remove progressive temporal downsampling by setting the stride of all temporal CNN layers to 1, while keeping the attention architecture unchanged, and apply a single temporal pooling operation only after all attention layers. As shown in Table IV, the performance of MaE spotting remains relatively stable under this single-scale setting due to the longer duration of MaEs. In contrast, ME spotting shows a substantial performance drop. This demonstrates that relying on a fixed temporal resolution is insufficient for effectively capturing subtle and short motion transients of MEs. Without multi-scale temporal modeling, these transient patterns become less salient during long-range temporal aggregation, as they may occupy only a small fraction of the input sequence. In contrast, progressive temporal downsampling facilitates hierarchical temporal abstraction, which better integrates discriminative cues for expressions of varying durations.
An example is shown in Fig. 6, where, without multi-scale temporal modeling, attention remains diffused across the full temporal resolution, while progressive temporal downsampling gradually concentrates attention on coarse temporal regions corresponding to the ME, thereby enabling the detection of the ME apex frame. Both quantitative and qualitative results indicate that proper multi-scale temporal modeling is critical for accurate ME spotting.
TABLE IV: Results of the ablation study on the multi-scale temporal modeling.
SAMM-LV CAS(ME)2
Temporal downsampling MaE ME Overall MaE ME Overall
Average pooling 0.4286 0.3755 0.4139 0.4917 0.1515 0.4609
Max pooling 0.4349 0.3113 0.4009 0.5110 0.0345 0.4713
Multi-scale (Ours) 0.4447 0.4281 0.4401 0.5061 0.2817 0.4841
@@ 4.4.1 Multiple model architectures
TABLE V: Comparison of multiple model architectures on SAMM-LV and CAS(ME)2.
@@
We validate and compare our proposed multiple multi-scale spatio-temporal model architectures on the SAMM and CAS(ME)2 datasets, and the experimental results are shown in Table V. With the exception of Paral-ST, our models use the same number of parameters for a fair comparison.
@@
TABLE VI: Comparison of different temporal window sizes (receptive fields) on SAMM-LV.
Window Size / MaE ME Overall
@@
4.4.2 Temporal window size
We further explore how much temporal information is needed to accurately evaluate one frame. Specifically, we compare different temporal window sizes (i.e., temporal receptive field) for the CSW-MRO on the SAMM-LV dataset, and the experimental results are presented in Table VI. The results indicate that the optimal performance is obtained when the temporal window size is set to 17, which demonstrates the effectiveness of our core idea: perceiving complete MEs and distinguishing between MaEs and MEs. Specifically, the temporal boundary for distinguishing MaEs and MEs is 0.5 seconds, which corresponds to 15 frames when the frame rate is set to 30 fps. A receptive field of 17 frames proves sufficient to perceive a complete ME and amplify the motion information that exists in MEs. When the model can accurately spot MEs, windows that exhibit greater intensity of motion, or windows that exhibit a continuous expression and are hard to perceive a trend of this expression occurring or ceasing within a short period, are more likely to contain MaEs.
TABLE VII: Results of the ablation study on supervised contrastive learning and the choice of hyperparameter.
@@ 4.4.2 Temporal window size
Figure 8: Visualization analysis of supervised contrastive learning. (a) PCA distributions of certain MaE and ME frames: left, without supervised contrastive learning; right, with supervised contrastive learning. (b) PCA distributions of certain ME and neutral frames: left, without supervised contrastive learning; right, with supervised contrastive learning.
Figure 9: Qualitative analysis of SpotFormer. The highlighted ROIs indicate the facial regions most relevant to expression dynamics and most influential to the model’s predictions.
4.4.3 Supervised contrastive loss and hyperparameter $\lambda$
Table VII shows the experimental results of introducing supervised contrastive learning into our model and the impact of the hyperparameter $\lambda$ in Eq. (5), which is set to balance classification and supervised contrastive learning. We observe that the introduction of the supervised contrastive loss enables our model to better recognize the classification boundaries between different types of expressions, resulting in an improvement of 6.0%/16.3% on the SAMM-LV and CAS(ME)2 datasets. The optimal value for $\lambda$ is 0.005, increasing $\lambda$ beyond this value starts to impact standard classification, leading to a decrease in spotting accuracy.
Fig. 8 shows the qualitative comparison between scenarios with and without supervised contrastive learning. In the inference phase, we randomly sampled frames and applied principal component analysis (PCA) [3] to analyze the distribution of various expression classes. Subsequently, we marked each frame with its corresponding ground-truth expression label. In Fig. 8 (a), we examine the PCA distribution of specific MaE frames and ME frames with and without supervised contrastive learning. Moving to Fig. 8 (b), we compare the PCA distribution of specific neutral frames and ME frames with and without supervised contrastive learning. The results indicate that, in the absence of supervised contrastive learning, our model may face challenges distinguishing certain MaE frames from ME frames due to subtle differences. This leads to mixed distributions and misclassification. Additionally, neutral frames may be mistakenly classified as ME frames, as the noise existing in the optical flow introduces slight fluctuations that the model may treat as ME motion information, resulting in incorrect ME proposal detection. With the introduction of supervised contrastive learning into our model, the domain discrepancy increases, thus significantly improving the accuracy of expression spotting.
@@ 4.4.3 Supervised contrastive loss and hyperparameter $\lambda$
TABLE VIII: Detailed spotting results of the proposed method on SAMM-LV and CAS(ME)2.
@@
4.5 Qualitative analysis
To better understand the interpretability of SpotFormer, we visualize the ROI-level spatial attention maps of two representative samples, which indicate how strongly each facial region contributes to the model’s prediction at each time point. Each sample includes the attention maps of the onset and apex frames, together with their corresponding raw images. As shown in Fig. 9, the highlighted ROIs correspond well to the facial areas involved in expression-related muscle activation. These visualizations demonstrate that SpotFormer can effectively localize the facial regions most relevant to expression dynamics, confirming its effectiveness in FES. The results further highlight the advantage of ROI-based modeling over methods that directly compress spatial information into a single feature vector, which may lose fine-grained local cues and become easily influenced by irrelevant motion noise.
@@ 4.6 Detailed results
Table VIII shows the detailed results on the SAMM-LV and CAS(ME)2 datasets. Our method achieves an overall precision rate of 0.4341/0.4781 and an overall recall rate of 0.4462/0.4902 on SAMM and CAS(ME)2, respectively. Specifically, our method achieves a recall rate of approximately 0.5 in MaE spotting and 0.4 in ME spotting. The exception is for ME spotting on the CAS(ME)2 dataset. The results indicate that it achieves a high precision rate while a low recall rate. The primary reason is that the scarcity and imbalance of ME data in the CAS(ME)2 dataset restrict the model’s ability to learn sufficiently discriminative representations for subtle ME patterns. CAS(ME)2 contains only 57 ME clips, and some samples involve only eye blinking—a motion generally considered irrelevant for FES. Due to the weak ME representations learned from such limited data, several true ME apex frames receive relatively low apex scores. Since our post-processing requires the apex probability to exceed a strict, dataset-independent threshold (0.5), these low-scoring true apex frames are filtered out, even though their scores remain higher than those of typical neutral frames. This results in missed detections and thus lower recall. Reducing the threshold specifically for ME spotting on CAS(ME)2 increases TPs but also introduces more FPs, indicating that the core issue lies in insufficient ME representation learning rather than threshold design.
We also found that, on both datasets, the precision is lower than the recall in MaE spotting, while the opposite is observed in ME spotting. This is mainly due to data imbalance and the specific characteristics of different expression classes in the FES datasets: MaEs are more frequent and generally exhibit higher intensity than MEs. This leads the model to focus more on MaEs and generate higher output probabilities for MaE spotting. As a result, more MaEs—including both TPs and FPs—are detected. In contrast, the model generates lower probabilities for ME spotting, resulting in fewer expression proposals but also fewer FPs.
@@
In summary, our proposed model can achieve a more advanced and balanced performance on both datasets, demonstrating the great generalization ability of our model. In addition, our method requires the detection of facial landmarks and the calculation of optical flows to form a head pose-irrelevant graph, making it more robust to variations such as random head angles. These merits demonstrate the potential of our method for practical applications (a cross-dataset evaluation on unseen data is provided in Section 4.7).
@@
TABLE IX: Results of cross-dataset validation.
SAMM-LV CAS(ME)2
Train Test MaE ME Overall MaE ME Overall
SAMM CAS(ME)2 - - - 0.3497 0.2000 0.3243
CAS(ME)2 SAMM 0.3390 0.1704 0.3008 - - -
4.7 Cross-dataset validation
To evaluate the robustness and generalization ability of our framework, we conduct cross-dataset validation, where the model is trained on one dataset and directly evaluated on the other without any fine-tuning. Specifically, we train the model on SAMM and evaluate it on CAS(ME)2, and vice versa. This setting simulates a real-world application scenario, as the target-domain distribution differs substantially from the source domain.
As shown in Table IX, our method achieves strong performance in both cross-dataset directions and consistently outperforms several existing methods, even though it is trained on a completely different dataset. This demonstrates the effectiveness of the proposed CSW-MRO feature and the ability of SpotFormer to capture transferable spatio-temporal cues rather than overfitting to dataset-specific patterns. Notably, despite the limited number of ground-truth ME samples in CAS(ME)2, the model still achieves an ME spotting F1-score of 0.1704 when trained on CAS(ME)2 and evaluated on SAMM, outperforming multiple prior methods. These results confirm the robustness of our framework under significant domain shift and highlight its potential applicability in real-world scenarios where the model must operate on unseen data.